# Imprints of interaction processes in the globular cluster system of NGC 3640

### v1

@@ Abstract
We present a wide-field study of the globular cluster systems (GCS) of the elliptical galaxy NGC 3640 and its companion NGC 3641, based on observations from Gemini Multi-Object Spectrograph/Gemini. NGC 3640 is a shell galaxy which presents a complex morphology, which previous studies have indicated as the sign of a recent ‘dry’ merger, although whether its nearest neighbour could have had an influence in these substructures remains an open question. In this work, we trace the spatial distribution of the globular clusters (GCs) as well as their colour distribution, finding a potential bridge of red GCs that connects NGC 3640 to its less massive companion, and signs that the blue GCs were spatially disturbed by the event that created the shells.
@@ 1 Introduction
Mergers are essential parts of the current $\Lambda$ Cold Dark Matter paradigm, in which dark matter haloes and galaxies are mainly formed through hierarchical assembly (Peebles, 1982; Blumenthal et al., 1984; Davis et al., 1985). In particular, the early evolution of early-type galaxies (ETGs) is dominated by major mergers involving large amounts of gas (Naab et al., 2007). Once they become quiescent, they continue to grow their mass and size through minor dry mergers, with little to no star formation (Naab et al., 2009, e.g.). Aside from being the driving forces behind the mass growth of a galaxy, mergers shape the morphology of galaxies (Hopkins et al., 2010; Kannan et al., 2015, e.g), feed central supermassive black holes, and potentially trigger central starbursts (Ellison et al., 2011; Satyapal et al., 2014). As such, they are important components of any galaxy formation model.
@@
Globular clusters (GCs) located in the halo of galaxies are helpful tools when it comes to tracing the occurrence of recent mergers since they are bright and they are connected to the underlying stellar population, acting as fossil records of the evolutionary history of their host galaxy and carrying in their properties valuable information about its accretion events. In regions where the stellar halo is too faint, GCs have been used as tracers of stellar streams and tidal tails (Napolitano et al., 2022), and when structures can be detected, the colours and positions of the GCs connected to them can shed further light on their origin (Lim et al., 2017; D’Abrusco et al., 2022, e.g). Since GCs are sparse even in massive galaxies and tidal features are hard to detect, studying this connection requires deep, wide-field observations. In nearby ETGs, rich underlying substructure is related with GC systems with unusual colour and luminosity distributions (e.g. Sesto et al., 2016; Bassino & Caso, 2017), and the presence of young GCs (e.g. Strader et al., 2004; Woodley et al., 2010).
@@ 2.1 Observations
The observations for this work were taken utilizing the Gemini Multi-Object Spectrograph Camera (GMOS) on Gemini North (Program GN-2016A-Q-69, PI: L. Bassino) on February 10th, 2016, in imaging mode. The field of view of GMOS is of $5.5\,\rm{arcmin}\times 5.5\,\rm{arcmin}$, and we used a binning of $2\times 2$, which results in a resolution of $0.146\,\rm{arcsecpix}^{-1}$. The seeing for these observations was $\approx 0.75$ across all filters. We observed two fields as shown in Figure 1, one containing both galaxies and an adjacent one with which we aimed to cover the entire extension of the GCS and calculate the background contamination. We also observed the Landolt field of standard stars SA 104 (Landolt, 1992), using the magnitudes obtained in the GMOS filters by Jørgensen (2009).
Figure 2: From top to bottom, the error in magnitude, $\chi^{2}$ and the sharpness value obtained from PSF photometry for the $g^{\prime}$ filter for reference. Solid lines indicate the chosen limits, with the limit in magnitude applied to the selections in the following tests.
@@ 2.2 Data reduction and point-source selection
The first catalog of globular cluster candidates present in the field was built using Source Extractor (Bertin & Arnouts, 1996). We used two filters, Gaussian and Mexhat ones, since the latter has proven to be more effective in crowded regions. Combining the outcome of both, we obtained approximately 4000 objects. Considering GCs appear as unresolved point-like sources at this distance, we use the classifier $CLASS\_STAR$, which estimates the probability of the detected source being point-like. In this case, we eliminate from the sample objects with $CLASS\_STAR>0.5$ as to clear it of extended objects, mainly background galaxies and saturated stars.
We obtained aperture photometry for these objects using the tasks in the DAOPHOT package in IRAF. After selecting $\sim 30$ of the brightest objects in each filter, for each field correspondingly, we built the spatially variable corresponding point-spread function (PSF) and obtained PSF photometry of our sample with the task ALLSTAR.
Figure 3: Completeness fraction in the $i^{\prime}$ filter for three concentric annular regions, with the corresponding fit. Circles and a solid line indicate the inner region; triangles and a dashed line, the intermediate region; and squares and a dotted line, the outer region.
Besides calculating instrumental magnitudes with their errors, this task provides us with statistical parameters that characterize the goodness of the fit. In particular, it calculates the sharpness and the chi-squared. In the completeness test described in the following section, we obtained the distribution of parameters for a large number of artificial sources, and selected constraints for these parameters, as shown in Figure 2. After applying this selection, we obtain a total of $1576$ point-like sources.
@@ 2.3 Completeness test
In the dim end of the magnitudes range, we estimate the limiting value by performing a completeness test. We build a catalog of 32000 objects, with magnitudes assigned so that they cover the range of magnitudes and colours expected for old GCs, and positions that cover both fields, distributed evenly over 200 copies of the original image. Using the PSF obtained before, we add them as fake stars to our raw data. Then, we proceed to perform the same data reduction process as before, detecting point-like sources and performing PSF photometry in the same manner as for our GC candidates. This allows us to obtain the set of statistical parameters that return the highest fraction of sources while still providing a good fit.
We divide the area into three concentric annular regions to take into consideration any radial variations in the completeness, and we estimate the fraction of recovered sources per magnitude bin for each of these regions. Then, we fit a function to the completeness fraction estimated for each region, which will be used later on to apply a more detailed completeness correction (Figure 3). This is particularly relevant in the inner region, in which the intrinsic brightness of the galaxy causes the completeness fraction to be lower than in other areas of the image. Finally, we determine the overall limit as the magnitude at which this fraction is of $\sim 80\%$ for the intermediate and outer regions, which in this case is $i^{\prime}_{0}=24.7$.
@@ 2.5 GC candidate selection
In Figure 4, we show the colour-colour diagrams of the entire sample of point-like sources detected in the field. We use colour limits to identify GC candidates, following Bassino & Caso (2017), in the three classic colour combinations. The limits in $(g^{\prime}-i^{\prime})$ are shown in the colour-magnitude diagram (Figure 6), where we add the limit obtained from the completeness test, corresponding to $80\%$, in the dim end of the magnitudes axis. The colour limits correspond to $1.50.4<(g^{\prime}-i^{\prime})_{0}<$ 1.5, $1.20.3<(g^{\prime}-r^{\prime})_{0}<$ 1.2 and $-0.05<(r^{\prime}-i^{\prime})_{0}<0.5$. In addition, we limit the error in magnitude to $<0.15$, and we exclude objects brighter than $M_{i}=21.5$, which at the estimated distance for NGC 3640 corresponds to $M_{V}=-11$, which is the limit at which the probability of finding GCs becomes very low (Harris et al., 2014). Objects brighter than this limit that fall in this colour range are more likely to be ultra-compact dwarfs, and as such, we treat them as potential contaminants.
Figure 5: Spatial distribution of all GC candidates across both fields. Orange dots and green diamonds indicate sources attributed to NGC 3640 and NGC 3641 respectively. The centre of each galaxy is shown in large symbols with thick edges.
Since we cannot estimate radial velocity measurements from photometry alone, initially we separate GC candidates according to their spatial distribution, as shown in Figure 5. The relation between the stellar mass of the galaxy and the extension of the GCS shown in Equation 9 in Caso et al. (2019) results in a value of $21\,kpc$ for the extension of the GCS of NGC 3641. However, the fact that NGC 3641 is a compact elliptical means that the estimated extension of the GCS that corresponds to the effective radius of the galaxy is far smaller. Since existing scaling relations are inconclusive for cE galaxies, we initially attribute all GC candidates within a radius of $75^{\prime\prime}$ of the center of NGC 3641 to its GCS, and the rest to NGC 3640, though this separation is by no means conclusive.
@@ 3.1 Colour-magnitude diagram
In Figure 6 we show the colour-magnitude diagram of the point-like sources in the field and the GCs associated with each galaxy. NGC 3641 presents intermediate colours in comparison with NGC 3640, covering a smaller range in general. In particular, NGC 3641 almost lacks blue GCs. It is reasonable to expect those to have been accreted first if these galaxies have interacted.
NGC 3640 covers a range typical of GCS in massive ETGs, and towards the brighter end, two subpopulations are evident.
@@ 3.2 Radial distribution
In Figure 7 we show the radial distribution for the GC candidates across both fields attributed to NGC 3640. We estimate the background level considering the flattening of the distribution in the outer regions, at $r_{gal}=3.7^{\prime}$, obtaining a contamination correction of $8\,\textrm{GC}\,\textrm{arcmin}^{-2}$ which was applied throughout the analysis.
@@
$n(r)=1.88\left(1+\left(\frac{r}{1.16}\right)^{2}\right)^{-2.06}$ (2)
Extrapolating this function to the value at which it reaches $20\%$ of the background level, we calculate a total extension of the system of $31\,{\rm kpc}$ ($\approx 4^{\prime}$).
Integrating the profile up to this point, we obtain the total population up to our completeness level, resulting in $\approx 310\,$GCs.
In addition, in Figure 8, we show the radial distribution for GC candidates split into red and blue subpopulations, following the separation described in the following Section. We fit Hubble-Reynolds laws to both profiles, obtaining core radii of $0.72$ for blue GCs and $1.8$ for red GCs. The red subpopulation appears heavily concentrated towards the centre of the galaxy, whereas the blue one extends to the outer region in a steeper manner. These behaviours are consistent with what is expected if we consider red GCs to be mostly formed in-situ, and blue GCs to come from accretion processes that increase their density in the outer halo.
Figure 7: Radial distribution for GC candidates attributed to NGC 3640. Crosses describe the raw sample, while orange circles show the background corrected distribution, with a grey, larger circle indicating those considered for the fit. The dash-dotted line shows the level at which $20\%$ of the background density is reached, and the dashed curve, the Hubble-Reynolds law fit.

### v2

@@ Abstract
We present a wide-field study of the globular cluster systems (GCS) of the elliptical galaxy NGC 3640 and its companion NGC 3641, based on observations from Gemini Multi-Object Spectrograph/Gemini using the $g^{\prime}r^{\prime}i^{\prime}$ filters. NGC 3640 is a shell galaxy which presents a complex morphology, which previous studies have indicated as the sign of a recent ‘dry’ merger, although whether its nearest neighbour could have had an influence in these substructures remains an open question. In this work, we trace the spatial distribution of the globular clusters (GCs) as well as their colour distribution, finding a potential bridge of red (metal-rich) GCs that connects NGC 3640 to its less massive companion, and signs that the blue (metal-poor) GCs were spatially disturbed by the event that created the shells. While NGC 3640 presents a typical bimodal colour distribution, the GCS attributed to NGC 3641 appears unimodal, with most GCs presenting an intermediate colour in comparison to those closer to its companion.
@@ 1 Introduction
Mergers are essential parts of the current $\Lambda$ Cold Dark Matter paradigm, in which dark matter haloes and galaxies are mainly formed through hierarchical assembly (Peebles, 1982; Blumenthal et al., 1984; Davis et al., 1985). In particular, the early evolution of early-type galaxies (ETGs) is dominated by major mergers involving large amounts of gas (Naab et al., 2007). Once they become quiescent, they continue to grow their mass and size through minor dry mergers, with little to no star formation (e.g Naab et al., 2009). Aside from being the driving forces behind the mass growth of a galaxy, mergers shape the morphology of galaxies (Hopkins et al., 2010; Kannan et al., 2015, e.g), feed central supermassive black holes, and potentially trigger central starbursts (Ellison et al., 2011; Satyapal et al., 2014). As such, they are important components of any galaxy formation model.
@@
Globular clusters (GCs) are considered fossil records of the stellar formation of their host galaxy due to their old ages, $10-13$ Gyr, (e.g. Strader et al., 2005; Tonini, 2013; Fahrion et al., 2020). They are also tracers of the evolutionary history of the galaxy, since they are survivors of its accretion events and disruption. GCs are almost ubiquitous in galaxies over $10^{9}\,M_{\odot}$ (Harris et al., 2015), and the properties of these globular cluster systems (GCS) have been linked to those of the halo mass (Hudson et al., 2014; Forbes et al., 2018; El-Badry et al., 2019) and to the environment in which the galaxy is located (De Bórtoli et al., 2022). These scaling relations are widely accepted and further the use of GCS as tracers of the merger histories of galaxies in the Local Universe (e.g Sesto et al., 2018; Kruijssen et al., 2020). In addition, GCs are intrinsically bright objects, which makes them detectable at distances of up to hundreds of Mpc (e.g Alamo-Martínez et al., 2013; Harris & Reina-Campos, 2024). In regions where the stellar halo is too faint, GCs have been used as tracers of stellar streams and tidal tails (Napolitano et al., 2022), and when structures can be detected, the colours and positions of the GCs connected to them can shed further light on their origin (Lim et al., 2017; D’Abrusco et al., 2022, e.g). Since GCs are sparse even in massive galaxies and tidal features are hard to detect, studying this connection requires deep, wide-field observations. In nearby ETGs, rich underlying substructure is related with GC systems with unusual colour and luminosity distributions (e.g. Sesto et al., 2016; Bassino & Caso, 2017), and the presence of young GCs (e.g. Strader et al., 2004; Woodley et al., 2010).
@@
Galaxy RA1 Dec1 Distance2 $M_{K}^{3}$ $V/\sigma_{e}^{3}$ $\rm{r}_{\rm{eff}}^{3}$
NGC 3640 11h21m06.8s +03d14m05s 26.9 Mpc -24.60 mag 0.370 $30.9\,\rm{arcsec}\,(1.9\,kpc)$
NGC 3641 11h21m08.8s +03d11m41s 26.9 Mpc -21.85 mag 0.331 $9.33\,\rm{arcsec}\,(1.2\,kpc)$
Table 1: Properties of both galaxies obtained from 1 NED, 2 SBF (Tully et al., 2013), and 3 Atlas3D Survey.
@@ 2.1 Observations
The observations for this work were taken utilizing the Gemini Multi-Object Spectrograph Camera (GMOS) on Gemini North (Program GN-2016A-Q-69, PI: L. Bassino) on February 10th, 2016, in imaging mode using the g’r’i’ filters with exposition times as listed in Table 2. The field of view of GMOS is of $5.5\,\rm{arcmin}\times 5.5\,\rm{arcmin}$, and we used a binning of $2\times 2$, which results in a resolution of $0.146\,\rm{arcsec\,pix}^{-1}$. The seeing for these observations was $\approx 0.75\,$arcsec across all filters. We observed two fields as shown in Figure 1, one containing both galaxies and an adjacent one with which we aimed to cover to cover most of the extension of the GCS. Based on the stellar mass of NGC 3640, $1.5\times 10^{11}\,{\rm{M}_{\odot}}$ (Karachentsev et al., 2015) we estimate the full extension of the GCS to be $r_{L}\approx 100\,\rm{kpc}$, i. e. $\sim 12$ arcmin according to De Bórtoli et al. (2022), although this is an upper limit, considering the environmental dependence suggested in that article for central galaxies.
We also observed the Landolt field of standard stars SA 104 (Landolt, 1992), using the magnitudes obtained in the GMOS filters by Jørgensen (2009).
Figure 2: From top to bottom, the error in magnitude, $\chi^{2}$ and the sharpness value obtained from PSF photometry for the $i^{\prime}_{0}$ filter for reference, although the selection was carried out in all filters. Solid lines indicate the chosen limits, with the limit in magnitude applied to the selections in the following tests.
@@ 2.2 Data reduction and point-source selection
The first catalog of objects present in the field was built using Source Extractor (Bertin & Arnouts, 1996). The run was performed twice using different filters each time, first a Gaussian filter which is considered best for detecting faint objects, and then a ‘mexhat’ filter which is recommended for crowded fields such as the region closest to the galaxy. The resulting catalogues were combined, ensuring an efficient recovery of sources. This process was done on the three available bands. Given that the resulting catalog in both filters for the $i^{\prime}$ band was the largest and contained all sources detected in $g^{\prime}$ and $r^{\prime}$, with a total of $\sim 6000$ sources, we chose it as our base catalog at this point. Typical effective radius of a GC is $r_{\rm{eff}}=3\,$pc (Jordán et al., 2007; Masters et al., 2010; Norris et al., 2014), which at the estimated distance to NGC 3640 corresponds to 0.030.030.03\,0.03arcsec. This is approximately two fifths of the value of the seeing of these observations, resulting in GCs appearing as unresolved point-like sources. We use then the classifier $CLASS\_STAR$ within Source Extractor, which estimates the probability of the detected source being point-like. In this case, we eliminate from the sample objects with $CLASS\_STAR>0.5$ to clear it of extended objects, mainly background galaxies and saturated stars ($CLASS\_STAR\sim 0$.)
We obtained aperture photometry for these point-like sources using the tasks in the DAOPHOT package in IRAF (Stetson, 1987). After selecting $\sim 30$ of the brightest objects in each filter, for each field correspondingly, we built the spatially variable corresponding point-spread function (PSF) and obtained PSF photometry of our sample in the three bands with the task ALLSTAR.
Figure 3: Completeness fraction in the $i^{\prime}$ filter for four concentric annular regions. From the innermost one outwards, they are indicated with circles and a solid line, triangles and a dashed line, squares and a dotted line, and crosses and a dot-dashed line. The magnitude is standardized as described in Section 2.4.
Besides calculating instrumental magnitudes with their errors, this task provides us with statistical parameters that characterize the goodness of the fit. In particular, it calculates the sharpness and the chi-squared for each source. In the next section, we describe the addition of artificial point-sources in the images of each band, to test the photometric completeness. Based on the parameters obtained from the PSF photometry for these artificial stars, we selected constraints for the sources from the science catalogue in each band, as shown in Figure 2. The constraints were analysed separately in each band, and resulted in similar enough values since the morphometric properties of the sources remained consistent across filters. Afterwards, we match the resulting catalogues and we obtain a total of $\sim 1300$ point-like sources.
@@ 2.3 Completeness test
In the dim end of the magnitudes range, we estimate the limiting value by performing a completeness test. This process was carried out in the three bands, although we focus on the result for the $i^{\prime}$ band since it was the deepest and thus the one used as a reference for the detection of sources. The detections in the $i^{\prime}$ band reached $1\,\rm{mag}$ past the expected turn-over magnitude, estimated to be $i^{\prime}_{0}=24.1$ as described in Section 3.4. In bands $g^{\prime}$ and $r^{\prime}$ there is a small difference in the faintest detection of approximately $0.2\,\rm{mag}$, which is expected since the integration times were chosen to achieve comparable depth.
We build a catalog of 32000 objects, following a typical GCS radial distribution and assigning magnitudes uniformly within the expected range for colors ($1.50.4<(g^{\prime}-i^{\prime})_{0}<$ 1.5 and $0.50<(r^{\prime}-i^{\prime})_{0}<$ 0.5) and magnitudes ($2616<i^{\prime}_{0}<$ 26) of GCs in massive galaxies (Harris et al., 2014). Using the PSF obtained before, we add them as fake stars to our raw data. Then, we proceed to perform the same data reduction process as before, detecting point-like sources and performing PSF photometry in the same manner as for our sources, and applying the same selection criteria described throughout this work, including the requirement that sources be detected in all three filters. This allows us to obtain the set of statistical parameters that return the highest fraction of sources while still providing a good fit. We matched the detected sources to the injected ones with a maximum spatial wandering of 0.70.70.7\,0.7arcsec, and performed a general check to ensure the detected magnitudes did not deviate more than $0.5$ mag from the originals.
We divide the area into four concentric annular regions limited by threee concentric circles centered in NGC 3640 to take into consideration any radial variations in the completeness, and we estimate the fraction of recovered sources per magnitude bin for each of these regions. Then, we fit the function defined in Equation 1 by Harris et al. (2009) to the completeness fraction estimated for each region, which will be used later on to apply a more detailed completeness correction (Figure 3). This is particularly relevant in the inner region, in which the intrinsic brightness of the galaxy causes the completeness fraction to be lower than in other areas of the image. It is to be noted that the outermost region corresponds to the adjacent field, where completeness is lower due to slightly different observing conditions. Finally, we determine the overall limit as the magnitude at which this fraction is of $\sim 80\%$ for the intermediate and outer regions, which in this case is $i^{\prime}_{0}=24.7$. This value is thus taken as our limit in the fainter end. To check whether there is a colour dependence in the completeness levels, we split the objects at an intermediate colour, $(g^{\prime}-i^{\prime})_{0}=0.9$ and analysed the completeness for the two resulting samples. At the limiting magnitude, objects bluer and redder than $(g^{\prime}-i^{\prime})_{0}=0.9$ were $2\%$ higher and $5\%$ lower than the whole sample, respectively, showing similar differences throughout the entire magnitude range, which renders the colour dependence neglectable in terms of this analysis.
@@ 2.4 Photometric calibration
Statistical tests from PSF fit $chi<0.9$
Statistical tests from PSF fit $||sharpness||<2$
Magnitude errors $m_{err}<0.15$
Completeness limit (faint end) $i^{\prime}_{0}=24.7$
Ultra-compact dwarf limit (bright end) $i^{\prime}_{0}=21.5$
Color limits $1.50.4<(g^{\prime}-i^{\prime})_{0}<$ 1.5
Color limits $1.20.3<(g^{\prime}-r^{\prime})_{0}<$ 1.2
Color limits $-0.05<(r^{\prime}-i^{\prime})_{0}<0.5$
Table 4: Adopted selection criteria for the GC candidates. Both the statistical tests from PSF fit and the magnitude errors were applied across all bands.
@@ 2.5 GC candidate selection
In Figure 4, we show the colour-colour diagrams of the entire sample of point-like sources detected in the field. We use colour limits to identify GC candidates in the three classic colour combinations, applying the same cuts as in Bassino & Caso (2017). The limits in $(g^{\prime}-i^{\prime})$ are shown in the colour-magnitude diagram (Figure 6), where we add the limit obtained from the completeness test, corresponding to $80\%$, in the dim end of the magnitudes axis. Across all bands the limits correspond to $1.50.4<(g^{\prime}-i^{\prime})_{0}<$ 1.5, $1.20.3<(g^{\prime}-r^{\prime})_{0}<$ 1.2 and $-0.05<(r^{\prime}-i^{\prime})_{0}<0.5$ mag. In addition, we limit the error in magnitude to $<0.15$ mag in all filters. As shown in Harris et al. (2014) objects brighter than $M_{V}=-11$ found around massive ETGs are more likely to be ultra-compact dwarfs, and the probability of finding GCs becomes very low. As mentioned in the Introduction, the SBF distance to NGC 3640 is $26.9$ Mpc (i.e. (m-M)=32.15), resulting in a cut-off magnitude on the brighter end of ${i^{\prime}_{0}}=21.5$. After applying the criteria based on the statistical tests from the PSF fit to the catalog that resulted from matching the sources detected in all three bands, we performed a new selection using these cuts in colour and magnitudes. All selection criteria are summarized in Table 4. This is the final catalog used for the analysis of the GCS, consisting of 725 objects.
Figure 5: Spatial distribution of all GC candidates across both fields. Orange dots and green diamonds indicate sources attributed to NGC 3640 and NGC 3641 respectively. Purple circles indicate objects within the colour ranges used for GCs considered background sources. The centre of each galaxy is shown in large symbols with thick edges.
Since we cannot estimate radial velocity measurements from photometry alone, initially we separate GC candidates according to their spatial distribution, as shown in Figure 5. The relation between the stellar mass of the galaxy and the extension of the GCS shown in Equation 9 in Caso et al. (2019) results in a value of $\sim 21\,$kpc (i.e. $\sim 2.6$ arcmin at the assumed distance) for the extension of the GCS of NGC 3641, considering a stellar mass of $10^{10}\,{\rm M_{\odot}}$ (Miller et al., 2015). The extension according to the effective radius of the galaxy, however, is $\sim 6\,$kpc (i.e. $\sim 0.8\,$arcmin) following the relation by Kartha et al. (2014). Considering this and the distance to NGC 3640 which creates a limitation to how far we can extend the analysis, we initially attribute all GC candidates within a radius of $1.25$ arcmin of the center of NGC 3641 to its GCS, and the rest to NGC 3640, though this separation is by no means conclusive.
@@ 3.1 Colour-magnitude diagram
In Figure 6 we show the colour-magnitude diagram of the point-like sources in the field and the GCs associated with each galaxy. NGC 3641 presents intermediate colours in comparison with NGC 3640, covering a smaller range in general. In particular, NGC 3641 almost lacks blue GCs. If the galaxies have interacted, it is reasonable to expect blue GCs initially from NGC 3641 to have been accreted first by NGC 3640, considering that blue GCs have more extended distributions than red ones (e.g. Strader et al., 2011; Kluge et al., 2023). These accreted blue GCs may now be mixed with those of NGC 3640 or they could present a much more extended spatial distribution which is limited in this analysis by the hard boundary set to separate both systems.
NGC 3640 covers a colour range typical of GCS in massive ETGs, and towards the brighter end, two subpopulations are evident.
@@ 3.2 Radial distribution
In Figure 7 we show the radial distribution, corrected by completeness, for the GC candidates across both fields attributed to NGC 3640. We estimate the background level considering the flattening of the distribution in the outer regions, at $r_{\rm{gal}}=3.7\,$arcmin, obtaining a contamination correction of $8\pm 3\,\textrm{GC}\,\textrm{arcmin}^{-2}$ which was applied throughout the analysis.
@@
$n(r)[N\rm{arcmin}^{-2}]=1.88\left(1+\left(\frac{r[\rm{arcmin}]}{1.16}\right)^{2}\right)^{-2.06}$ (2)
Previous studies of GCS on wide fields of view have shown that adopting a projected density of 0.2-0.3 times the background density results in consistent measurements of the extension of the GCS (e.g. Bassino et al., 2006; Caso et al., 2017). We extrapolate then this function to the value at which it reaches $20\%$ of the background level, and calculate a total extension of the system of $31\,{\rm kpc}$ ($\approx 4\,$arcmin).
Integrating the profile up to this point, we obtain the total population up to our completeness level, resulting in $268\pm 9\,$GCs.
In addition, in Figure 8, we show the radial distribution for GC candidates split into red and blue subpopulations, following the separation described in the following Section. We fit Hubble-Reynolds laws to both profiles, obtaining core radii of $0.72$ arcmin for red GCs and $1.8$ arcmin for blue GCs. The red subpopulation appears heavily concentrated towards the centre of the galaxy, whereas the blue one extends to the outer region in a steeper manner. These behaviours are consistent with what is expected if we consider red GCs to be mostly formed in-situ, and blue GCs to come from accretion processes that increase their density in the outer halo (e.g. Reina-Campos et al., 2019; El-Badry et al., 2019; Harris & Reina-Campos, 2023).
Figure 7: Radial distribution for GC candidates attributed to NGC 3640. Crosses describe the raw sample, while orange circles show the background corrected distribution, with a grey, larger circle indicating those considered for the fit. The dashed line shows the level at which $20\%$ of the background density is reached, and the dash-dotted curve, the Hubble-Reynolds law fit.